Please help me with
The product of two number is 442. The sum of the numbers is 43.
x*n=442
x+n=43

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

The product of two number is 442. The sum of the numbers is 43.
x*n=442
x+n=43
--------
Solve the 2nd equation for x = get x=-n+43
Substitute that into the 1st equation to get:
(-n+43)n = 442
43n-n^2 = 442
n^2-43n+442=0
n = [43+-sqrt(43^2-4*442)]/2
n = [43+-9]/2
n = 52/2 = 26
Substitute that into x+n=43 to get x = 17
================
Cheers,
Stan H.

Answer by checkley71(8403)

X+Y=43
X=43-Y
X*Y=442
(43-Y)Y=442
43Y-Y^2=442
Y^2-43Y+442=0
(Y-26)(Y-17)=0
Y-26=0
Y=26 ANSWER.
Y-17=0
Y=17 ANSWER.
PROOF
17+26=43
43=43
17*26=442
442=442